### v1

@@ Acknowledgments
AH, BM, and NM acknowledge support from the Physics and Engineering Models (PEM) Materials project within the Advanced Simulation and Computing (ASC) Program at LANL (ASC-PEM-Materials Project). AM and NM also acknowledges funding from the LDRD-DR “Investigating How Material’s Interfaces and Dislocations Affect Strength” (iMIDAS) under Grant No. 20210036DR (Abigail Hunter, PI and Saryu J. Fensin, Co-PI) and for the initial MD calculations and methodology. BM also acknowledges Ian Chesser and Ian Wise for their constructive comments and discussions during the development of the work. LA-UR-25-26159.
@@
Basic scheme
@@
In the particular case of isotropic elasticity it may redefined as
Details of the coupled formulation
This sections aims to show some of the steps that were not shown in Section 2 to get Equations 11, 13 and 20.
Equation 11 is obtained by following these steps: From Equation 4 and applying the balance of linear momentum one gets:
rearranging the equation above and getting the total strain multiplied by the reference elastic modulus on the left hand side, one gets
using Green’s operator and the Fourier transform the solution for the total strain is
Equation 11 is obtained by replacing the equation of total strain presented above into the strain decomposition (Equation 2):
The Fourier transform of the elastic strain is used to get the equation for the elastic energy
replacing the elastic strain into the equation above, the elastic energy may be redefined as
multiplying the two elastic strains the elastic energy gives the same expression shown in Equation 13
@@
rearranging the terms and dividing by the volume
note that by definition the term . Equaton 20 is obtained using the definitions of averaged quantities presented by the end of Section 2 (, , , and ). For the case of applied strain and . If , the average strain is defined as
which is equivalent to the formulation for a homogeneous material.
Calculation of second order elastic constants
Phase field dislocation dynamics simulations
In the main text only three cases were presented, the remaining simulations are shown in the Supplementary Figures 2-15.
Averaged strain-gradient metrics calculated for classification in [Mishra et al.,, 2024] is plotted as a function of in Figures 16 and 17. These were computed using the small atomistic simulation cells used for GB width and elastic stiffness calculations. The average values were computed over a region around the GB corresponding to the GB width. These show that the are correlated with strain-gradient metrics. These GBs had the same macroscopic degrees of freedom and similar energy values to the GBs used in dislocation-GB interaction study Dang et al., [2025].
The data is more noisy when computed from large simulation cells as shown in Figures 18 and 19. In this case the averages were computed using all atoms in a region around the dislocation interaction site, defined by the GB width (normal to the GB plane) and 60 Å(normal to the tilt axis). As the larger atomistic simulation cells are less constrained, the GB undergoes relaxation resulting in differences in energy values compared to the small cells.

### v2

@@ Acknowledgments
AH, BM, and NM acknowledge support from the Physics and Engineering Models (PEM) Materials project within the Advanced Simulation and Computing (ASC) Program at LANL (ASC-PEM-Materials Project). AM and NM also acknowledges funding from the LDRD-DR “Investigating How Material’s Interfaces and Dislocations Affect Strength” (iMIDAS) under Grant No. 20210036DR (Abigail Hunter, PI and Saryu J. Fensin, Co-PI) and the LDRD-PRD project 20220814PRD4: “Grain Boundary Characterization from Diffractograms by Physics-Informed Machine Learning” for the initial MD calculations and methodology. BM also acknowledges Ian Chesser and Ian Wise for their constructive comments and discussions during the development of the work. LA-UR-25-26159 version 2.
Data availability
The data cannot be made publicly available upon publication because they are not available in a format that is sufficiently accessible or reusable by other researchers. The data that support the findings of this study are available upon reasonable request from the authors.
Conflicts of interest
The authors declare no conflicts of interest.
@@
1 Micromechanics solver: Basic scheme
@@
where is the wave number vector. In the particular case of isotropic elasticity it may redefined as
2 Details of the coupled formulation
This sections aims to show some of the steps that were not shown in Section 2 to get the derivative of the elastic energy density with respect to the order parameter and the macroscopic strain in Equation 16.
The elastic energy is computed in the real space using Equations 2, 8 and 10. The derivative of the elastic energy density with respect to the order parameter is computed as
with
Assuming that and , Equation S.4 is deifned as
Thus, Equation S.3 is redefined as
where the second term is solved in the Fourier space
In a compact form, it can be expressed as
with the interaction matrix defined as
@@ 2 Details of the coupled formulation
Rearranging the terms and dividing by the volume gives
Note that by definition the term . Equaton 16 is obtained using the definitions of averaged quantities presented by the end of Section 2 (, , , and ).
3 Phase field dislocation dynamics simulations
In the main text only three cases were presented, the remaining simulations are shown in Supplementary Figures 1-14.
Averaged strain-gradient metrics calculated for classification in [Mishra et al., 2024] are plotted as a function of in Figures 15 and 16. These were computed using the small atomistic simulation cells used for GB width and elastic stiffness calculations. The average values were computed over a region around the GB corresponding to the GB width. These show that the are correlated with strain-gradient metrics. These GBs had the same macroscopic degrees of freedom and similar energy values to the GBs used in dislocation-GB interaction study [Dang et al., 2025].
The data shows similar trends when computed from large simulation cells as shown in Figures 17 and 18. In this case the averages were computed using all atoms in a region around the dislocation interaction site, defined by the GB width (normal to the GB plane) and 60 Å(normal to the tilt axis). As the larger atomistic simulation cells are less constrained, the GB undergoes relaxation resulting in differences in energy values compared to the small cells.